OpenStudy (anonymous):

write the ratio as a ratio of whole numbers in lowest terms. $3.50 to $1.50

OpenStudy (anonymous):

3.50/1.50 = 7/3=7:3

OpenStudy (anonymous):

thanks! how's u do that?

OpenStudy (matt101):

You just multiply the numerator and denominator by the same factor - in this case 2.

OpenStudy (anonymous):

ok but i dont get how u got 7:3

OpenStudy (matt101):

The ratio can be expressed as a fraction (7/3), though sometimes people prefer to use the colon which means the same thing but only in the context of ratios. If you write 7:3, you say "7 to 3" so its obvious its a ratio, as oppose to saying "7 over 3" for a fraction which is more general. But it's just semantics haha...

OpenStudy (anonymous):

no thats not what I meant... it's ok thanks anyway

OpenStudy (matt101):

Well maybe you could be more specific with what you don't get? I don't mind giving you a hand :)

OpenStudy (anonymous):

thanks I hate ratios:( ugh soo annoying

OpenStudy (anonymous):

i have 12 problems to do for hw:/

OpenStudy (anonymous):

matt101 \[\frac{3.50/.50}{1.50/.50} =\frac{7}{3} \]

OpenStudy (matt101):

I said that...it's the same as multiplying by 2. NeeNee didn't get something else...

OpenStudy (anonymous):

it's you guys helped alot

OpenStudy (anonymous):

okay*